im studying for my end of course exam so anything on percent decrease/increase would be nice. thanks.
Answer by stanbon(75887)

im studying for my end of course exam so anything on percent decrease/increase would be nice. thanks
---
%Increase from $20 to $35??
---
1st figure out the QUANTITY INCREASE:
35-20 = 15
2nd divide the quantity increase by the initial price.
15/20 = 0.75 = 75% increase
====================================
% Decrease from $50 to $20???
1st figure out the QUANTITY DECREASE:
50-20 = 30
2nd divide the quantity decrease by the initial price
30/50 = 3/5 = 0.60 = 50%
=====================================
Hope this helps.
Cheers,
Stan H.